OpenStudy (karatechopper):

who has an Algebra 1 question

myininaya (myininaya):

4x=2 how do i solve?

OpenStudy (karatechopper):

divide by 4 on both sides

OpenStudy (karatechopper):

then simplify

myininaya (myininaya):

why can't i multiply both sides by 1/4 instead?

OpenStudy (karatechopper):

?????

myininaya (myininaya):

lol

myininaya (myininaya):

i'm jk is this consider an algebra1 question?

OpenStudy (karatechopper):

sure

OpenStudy (karatechopper):

it could be!!!

myininaya (myininaya):

what is the difference btw algebra 1 and 2?

OpenStudy (karatechopper):

i was thinking if anybody needed help tho....

OpenStudy (karatechopper):

uhhh algebra 2 comes after geometry?!?!

myininaya (myininaya):

you wanted to seem inviting

myininaya (myininaya):

algebra 2 comes after geometry?

OpenStudy (karatechopper):

yea i figured tht out

OpenStudy (karatechopper):

yep it does!!

myininaya (myininaya):

so you are talking about high school math

OpenStudy (karatechopper):

yea!!

myininaya (myininaya):

i remember this now i think it was: algebra 1 geometry algebra 2

OpenStudy (karatechopper):

yes it is!

myininaya (myininaya):

why don't you want a more challenging question? like a calculus problem?

OpenStudy (karatechopper):

b/c i am in 7th grade learning Algebra 1

myininaya (myininaya):

oh i see

OpenStudy (karatechopper):

and i only kno algebra 1

myininaya (myininaya):

ok how do i solve \[a^2+2ab+b^2+4a+4b=-4\] for a+b?

OpenStudy (karatechopper):

so the equation has to equal a+b?

OpenStudy (anonymous):

here is an algebra question. prove A5 is simple

OpenStudy (karatechopper):

and what???? satellite? i am only gonna focus on one question at a time.... and tht is myin's

myininaya (myininaya):

\[a^2+2ab+b^2=(a+b)^2\] \[4a+4b=4(a+b)\] \[(a+b)^2+4(a+b)=-4\] \[(a+b)^2+4(a+b)+4=0\]

OpenStudy (anonymous):

oh sorry. i will be quiet

OpenStudy (karatechopper):

now its ok cuz i dont get what the heck myin is sayin...

myininaya (myininaya):

lol

OpenStudy (karatechopper):

A5 IS A times 5

myininaya (myininaya):

asking you to solve that equation for a+b

OpenStudy (karatechopper):

ok now lemme get to work

myininaya (myininaya):

\[(a+b+2)^2=0\]

myininaya (myininaya):

\[a+b+2=0\] you can solve this for a+b right?

OpenStudy (karatechopper):

yes!!! bbut i am on the first equation hold on

myininaya (myininaya):

ok \[a+b=-2 :)\]

OpenStudy (karatechopper):

o wait i am confused!!!??? on first equation solve for me??? too many numbers...... the one u just said isa+b=-2

myininaya (myininaya):

\[a^2+2ab+b^2+4a+4b=-4\] \[(a+b)^2+4a+4b=-4\] \[(a+b)^2+4(a+b)=-4\] \[(a+b)^2+4(a+b)+4=0\] ---------------------------- remember how to factor \[u^2+4u+4\] \[(u+2)^2\] ---------------------------- \[(a+b+2)^2=0\] \[a+b+2=0 => a+b=-2\]

myininaya (myininaya):

do you guys factor in algebra 1? solve equations in algebra 1?

OpenStudy (karatechopper):

we havent factored yet..

myininaya (myininaya):

oh then thats the problem

myininaya (myininaya):

ok try satellite's question whatever it was

OpenStudy (karatechopper):

i already did

OpenStudy (anonymous):

is solve for x: \[2x-1=7\] and algebra 1 question? and if so how about \[ax+b-c\]?

OpenStudy (anonymous):

i meant \[ax+b=c\]

OpenStudy (karatechopper):

x=4 c-b/a=x

myininaya (myininaya):

ax=c-b x=(c-b)/a

OpenStudy (anonymous):

ok so i guess that takes care of that kind. how about \[9x+5=7x+1\] or \[ac+b=cx+d\]